Tomato weights and Fertilizer: Carl the farmer has three fields of tomatoes, on one he used no fertilizer, in another he used organic fertilizer, and the third he used a chemical fertilizer. He wants to see if there is a difference in the mean weights of tomatoes from the different fields. The sample data is given below. The second table gives the results from an ANOVA test. Carl claims there is a difference in the mean weight for all tomatoes between the different fertilizing methods.

Tomato-Weight in Grams

x
No Fertilizer 123   119     118     120     117     120     114     118     129     128     120.6  
Organic Fertilizer 112   127     138     133     140     114     126     134     123     144     129.1  
Chemical Fertilizer 115   141     143     134     129     134     135     129     113     148     132.1  

ANOVA Results

F P-value
4.040 0.0292

The Test: Complete the steps in testing the claim that there is a difference in the mean weight for all tomatoes between the different fertilizing methods.

(a) What is the null hypothesis for this test?

H0:  μ1 = μ2 = μ3.H0:  μ1μ2μ3.    H0: At least one of the population means is different from the others.H0:  μ3 > μ2 > μ1.


(b) What is the alternate hypothesis for this test?

H1: At least one of the population means is different from the others.H1:  μ1 = μ2 = μ3.    H1:  μ1μ2μ3.H1:  μ3 > μ2 > μ1.


(c) What is the conclusion regarding the null hypothesis at the 0.01 significance level?

reject H0fail to reject H0    


(d) Choose the appropriate concluding statement.

We have proven that all of the mean weights are the same.There is sufficient evidence to conclude that the mean weights are different.    There is not enough evidence to conclude that the mean weights are different.


(e) Does your conclusion change at the 0.10 significance level?

YesNo    
Answer #1

Step 1: State null and alternative hypotheses
Ho: μ1= μ2= μ3
H1: At least one mean is different from the others


Step 2: Find the critical value
k = 3
N = total number of data values from all groups = 33
d.f.N. = k -1 = 2
d.f.D. = N -1 = 30
α = 0.1
Critical Value = 2.488

Step 3: Calculate F Test Value
ΣX = Sum of all data values = 1326.6+1420.1+1453.1 = 4199.79
N = 33
Grand Mean = XGM = 4199.79/33 = 127.266

Calculate between-group variance, denoted by SB2:

SSB = 11(120.6 - 127.26)2+ 11(129.1 - 127.26)2+ 11(132.1 - 127.26)2

SSB = 782.83

S2B = 782.83 / 2 = 391.416

Calculate within-group variance, denoted by SW2:

SSW = (11- 1)(20.44)+ (11- 1)(103.09)+ (11- 1)(114.29)

SSW = 2378.20

N - k = 33 - 3 = 30

SW2 = SSW/(N - k) = 2378.20/30 = 79.27

Calculate F test value:

F = 391.41 / 79.273 = 4.9375

Step 4: Make Decision:
Compare F test value = 4.937557816836262 with Critical Value = 2.4887

The decision is to reject the null hypothesis since 4.9375 > 2.4887



Analysis of Variance Summary Table

Source    Sum of Squares    Degrees of Freedom (df)    Mean Square    F test value   
Between    SSB = 782.83 k - 1 = 2    MSB = 391.416   F = 4.9375
Within (error)    SSW = 2378.20 N - k = 30    MSW = 79.273      
Total    3161.03   32